OpenStudy (anonymous):

Which of the following is a solution to the equation y = 3x – 1? (4, 1) (2, 5) (4, 3) (0, –3)

OpenStudy (anonymous):

@Gabylovesyou

OpenStudy (gorv):

take (4,1) 1=3*4-1=12-1=11 1=11 false not a solution

OpenStudy (gorv):

similar check for other if both side comes equal that will be our solution

OpenStudy (anonymous):

so which one is it ?

OpenStudy (gorv):

that u try u will get the ans if u try